1. To prepare a buffer solution with a pH of 3.7. Which of the following conjugated pairs would be the best option?

A. HSO4- / SO42- Ka = 1.2 x10-2

B. HCN / CN- Ka = 4.0 x 10-10

C. HCHO2 / CHO2- Ka = 1.8 x 10-4

D. H2CO3 / HCO3 Ka = 4.2 x 10-7

E. HClO / ClO- Ka = 2.9 x 10-8
Answer #1

To prepare a buffer solution with a pH of 3.7. The pKa value of the weak acid should be near to 3.7.

pKa = -logKa

A. HSO4- / SO42- Ka = 1.2 x10-2

pKa = -log(1.2x10-2) = 1.92

This pKa is much lower than 3.7. So, this option is wrong.

B. HCN / CN- Ka = 4.0 x 10-10

pKa = -log(4.0x10-10) = 9.40

This pKa is much higher than 3.7. So, this option is also wrong.

C. HCHO2 / CHO2- Ka = 1.8 x 10-4

pKa = -log(1.8x10-4) = 3.74

This pKa is nearly equal to 3.7. So, this may be the best conjugated pair

D. H2CO3 / HCO3 Ka = 4.2 x 10-7

pKa = -log(4.2x10-7) = 6.34

This pKa is higher than that of 3.7. So, this option is wrong.

E. HClO / ClO- Ka = 2.9 x 10-8

pKa = -log(2.9x10-8) = 7.54

Thia pKa is much higher than that of 3.7. So, this option is also wrong.

Hence,

Answer: C

Let me know if you have any queries regarding this